OpenStudy (anonymous):

If Integral of f(x)d(t-a) with limits from (-infinity to + infinity) is f(a) where d is the delta function...what will be the case if limits are finite?

OpenStudy (anonymous):

will it be the same and limits will not affect the end result??

OpenStudy (anonymous):

\[\int_{-\infty}^{\infty} f(x) \delta(x-a) dx=f(a)\] ?

OpenStudy (anonymous):

yup @mukushla ...but for finite limits??

OpenStudy (turingtest):

\[\int_{a-\epsilon}^{a+\epsilon}f(t)\delta(t-a)=f(a)\]who says it needs to be infinite? just symmetrical

OpenStudy (anonymous):

well then we have \[\int_{a}^{\infty} f(x) dx=f(a)\]

OpenStudy (anonymous):

Max? do u know what he means?

OpenStudy (anonymous):

but if limits are like....2 to 8..then??

OpenStudy (anonymous):

what is the property of the delta function that will affect the integrand??

OpenStudy (turingtest):

2=5-3 8=5+3 epsilon =3 a=5

OpenStudy (turingtest):

\[\int_{5-3}^{5+3}f(t)\delta(t-a)dt=f(5)\]

OpenStudy (turingtest):

I suppose

OpenStudy (anonymous):

yup.....everything is repeating

OpenStudy (turingtest):

yes, lag+repeating post=awkward explanation

OpenStudy (turingtest):

erm, read above while I reload

OpenStudy (turingtest):

the integral has to be symmetrical

OpenStudy (anonymous):

let me give an example\[\int\limits_{-2}^{4}(2+t^2)\delta(t+1)dt\]...

OpenStudy (anonymous):

i have posted an example above..

OpenStudy (anonymous):

not necessary that it should be symmetrical

OpenStudy (anonymous):

i think when delta is a fraction such that its integrand denominator = 0 by substituting the limits ,this can also affect the results

OpenStudy (anonymous):

didnt understand!!!

OpenStudy (turingtest):

I think you were right above actually, the integral does not need to be symmetrical integration over any region containing the point x=a should give the same result the symmetry bit is needed to prove various properties

OpenStudy (turingtest):

in your integral a=-1 your range is (-2,4) and a=-1 is in that interval so it is jus t\( f(a)\)

OpenStudy (anonymous):

no....@TuringTest ...it will not be the answer.....i am trying to refer some books online.....lets see if i can get it...

OpenStudy (turingtest):

I was looking at this for online stuff http://tutorial.math.lamar.edu/Classes/DE/DiracDeltaFunction.aspx

OpenStudy (turingtest):

"Note that the integrals in the second and third property are actually true for any interval containing t=a, provided it’s not one of the endpoints" so it says

OpenStudy (turingtest):

I'm not getting much out of mathworld http://mathworld.wolfram.com/DeltaFunction.html

OpenStudy (fwizbang):

The delta function is the limit of some other set of functions(say Gaussians) that satisfies the criterion of zero width and unit integral in the limt., like \[\delta(x-a)= \lim_{\epsilon\rightarrow 0}\frac{1}{2\epsilon} if -\epsilon < x-a < \epsilon \] and 0 otherwise To use this in an integral that has a as an endpoint, you need to be very careful about how you go about taking the limits.

OpenStudy (anonymous):

Any points on either side will do.